OpenStudy (anonymous):

At what altitude above the earth's surface would the acceleration due to gravity be 4.9ms−2 ? Assume the mean radius of the earth is 6.4×106 metres and the acceleration due to gravity 9.8ms−2 on the surface of the earth

OpenStudy (shamim):

the formla will b\[g \prime =\frac{ GM }{ (R+h)^{2} }\]

OpenStudy (y2o2):

\[\large g= {G M_e \over r^2 } = 9.8\] \[\large g' = {G M_e \over (r + h)^2 } = 4.9\] \[\large 4.9(r+h)^2 = 9.8 r^2\] you can easily find h

OpenStudy (shamim):

\[g=\frac{ GM }{ R ^{2} }\]

OpenStudy (shamim):

now u hv to divide these 2 equation